# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
$s_{s_{0}}=_{d}(2s_{0})^{-1}\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right).$ (62)
Now that we have the explicit expression for the score vector $\mathbf{s}_{\bm{\eta}_{0}}$ in the canonical model $\mathcal{P}_{\bm{\eta},g}$ we can easily evaluate the related FIM $\mathbf{I}_{\bm{\eta}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\eta}_{0}}\mathbf{s}_{\bm{\eta}_{0}}^{T}\}$. Before computing FIM, we define two functionals of the true density generator $g_{0}$:
$\alpha(g_{0})\triangleq\frac{E\{\mathcal{Q}^{2}\varphi^{2}_{0}(\mathcal{Q})\}}{m(m+2)},$ (63)
$\beta(g_{0})\triangleq\frac{E\{\mathcal{Q}\varphi^{2}_{0}(\mathcal{Q})\}}{m}.$ (64)
@@
where  [^3] Note that, in the calculations, we used the equality $E\{\mathcal{Q}\varphi_{0}(\mathcal{Q})\}=m$. In fact, from (58), (33) and (30), we have $E\{\mathcal{Q}\varphi_{0}(\mathcal{Q})\}=-2\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2}dg=-2\delta_{m}^{-1}[q^{m/2}g(q)]_{0}^{\infty}+m\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2-1}g(q)dq=m$.
$\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\}=s_{0}^{-1}\beta(g_{0})\mathbf{V}_{S,0}^{-1},$ (66)
@@
$\begin{split}\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}&\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}\}=\\ &=\frac{1}{4}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\alpha(g_{0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})+\left(\alpha(g_{0})-1\right)\mathrm{vec}(\mathbf{I}_{m})\mathrm{vec}(\mathbf{I}_{m})^{T}\right]\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\right]^{T},\end{split}$ (69)
$\begin{split}I_{s_{0}}\triangleq E_{0}\{(s_{s_{0}})^{2}\}=\frac{m(m+2)\alpha(g_{0})-m^{2}}{4s_{0}^{2}},\end{split}$ (70)
$\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}s_{s_{0}}\}=\frac{(m+2)\alpha(g_{0})-m}{4s_{0}}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1}).$ (71)
We note that this result is in line with the expression already provided in [10, eqs. (6) and (7)].
We conclude this section by noticing that the score vector $\mathbf{s}_{\bm{\zeta}_{0}}$ for the nuisance parameters $\bm{\zeta}_{0}\in\Delta$, characterizing de density generator $g_{\bm{\zeta}}$ in $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ in (48), can be straightforwardly obtained from (49) as:
$\mathbf{s}_{\bm{\zeta}_{0}}=_{d}\frac{1}{g_{\bm{\zeta}_{0}}(\mathcal{Q})}\bm{\nabla}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q}).$ (72)
Now that we have clearly defined the statistical models and sub-models to be analyzed along with their corresponding score vectors, we can move on to analyze their geometric relationships and the consequent impact in terms of statistical efficiency.
IV The geometry of $\mathcal{P}_{\bm{\eta},g}$ and some results on parametric and semiparametric information matrices
In this section, we will specialise Lemmas II.1 and II.2, introduced in Sect. II, to the study of the elliptical distributions. The semiparametric model from which we start is the “canonical” one, i.e. $\mathcal{P}_{\bm{\eta},g}$ given in (40), corresponding to (c.t.) $\mathcal{P}$ in (5), along with its parametric submodels $\mathcal{P}_{\bm{\gamma}}$ in (44) (c.t. $\mathcal{P}_{1}$ in (6)), $\mathcal{P}_{s}$ in (45) (c.t. $\mathcal{P}_{2}$ in (7)), $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) and the non-parametric model $\mathcal{P}_{g}$ in (47) (c.t. $\mathcal{P}_{3}$ in (9)). Moreover, we will investigate the link between the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$, introduced in (48), with the above-mentioned models. The key result that allows us to study the impact of finite and infinite dimensional nuisance parameters on the semiparametric efficiency in $\mathcal{P}_{\bm{\eta},g}$ for the parameter vector of interest $\bm{\gamma}_{0}$ is Lemma II.2. To specialize this lemma to our context, we need to draw the following connections:
  • the finite-dimensional vector of the parameters of interest is $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ and the related score is
@@
    where $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ are given in (56) and (61), respectively.
  • the finite-dimensional nuisance parameter is $s_{0}$ and the related score $s_{s_{0}}$ is the one given in (62). As a consequence, from (14), the orthogonal projection onto the finite dimensional nuisance tangent space $\mathcal{H}\supseteq\mathcal{T}_{s_{0}}=\{a\cdot s_{s_{0}},a\in\mathbb{R}\}$ (c.t. $\mathcal{T}_{2}$ in (12)) is given by:
    $\Pi(h|\mathcal{T}_{s_{0}})=E_{0}\{hs_{s_{0}}\}I_{s_{0}}^{-1}s_{s_{0}},\;h\in\mathcal{H},$ (74)
@@
  • The infinite-dimensional nuisance tangent space $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$) is given in Appendix .7 as:
    $\begin{split}\mathcal{T}_{g_{0}}&=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(s_{0}^{-1}(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{V}_{S,0}^{-1}(\mathbf{x}-\bm{\mu}_{0})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h\;\text{is $\sigma(\mathcal{Q})$-measurable}\right\},\end{split}$ (75)
    where we used the definition of the transformation $Q_{\bm{\mu},\bm{\Sigma}}(\mathbf{x})$ in (32) and where $\sigma(\mathcal{Q})\subset\mathfrak{B}(\mathcal{X})$ is the sub-$\sigma$-algebra generated by the random variable $\mathcal{Q}$ in (32). Moreover, from e.g. [25, Ch. 23, Def. 4], we have that the orthogonal projection of a generic element $h\in\mathcal{H}$ onto $\mathcal{T}_{g_{0}}$ can be obtained as
    $\Pi(h|\mathcal{T}_{g_{0}})=E\{h|\mathcal{Q}\},\;\forall h\in\mathcal{H}.$ (76)
    Consequently, $\mathcal{H}\ni h\perp\mathcal{T}_{g_{0}}$ iff $E\{h|\mathcal{Q}\}=0$.
We are now ready to prove the following (surprising) result:
@@
The efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\eta},g}$ in (40) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$ in (46). Moreover, if the scale functional $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then the parametric FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ in the parametric model $\mathcal{P}_{\bm{\gamma}}$ in (44) is equal to the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for $\mathcal{P}_{\bm{\eta},g}$.
This surprising result had already been discovered in [10, 9] using a semiparametric generalization of Le Cam’s theory [20]. The geometric approach, based solely on the tools discussed in Sect. II, provides an alternative proof that may be more accessible to a wider audience, as it requires only a basic background in Hilbert spaces.
@@
From Lemma II.2, to prove the first part of Proposition IV.1, we need to show that
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\triangleq\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\perp\mathcal{T}_{g_{0}},$ (77)
which, by (73) and (76), is equivalent verifying that:
$E\{[\mathbf{s}_{\bm{\mu}_{0}}-\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{s_{0}})]|\mathcal{Q}\}=\mathbf{0}.$ (78)
$E\{[\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})]|\mathcal{Q}\}=\mathbf{0},$ (79)
where the score vectors $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ are given in (56) and (61), respectively. Let us start with the condition (78) on the location parameter $\bm{\mu}_{0}$. From the definition of the projection operator onto $\mathcal{T}_{s_{0}}$, i.e. $\Pi(\cdot|\mathcal{T}_{s_{0}})$, given in (74) and from the fact that $\mathbf{I}_{\bm{\mu}_{0},s_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}s_{s_{0}}\}=\mathbf{0}$ as indicated in (68), we immediately have that:
$\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{s_{0}})=\mathbf{I}_{\bm{\mu}_{0},s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=\mathbf{0}.$ (80)
Moreover, from the expression of the score $\mathbf{s}_{\bm{\mu}_{0}}$ given by (56), the independence between $\mathcal{Q}$ and $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the property $E\{\mathbf{u}\}=\mathbf{0}$, we have:
$E\{\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{Q}\}=\mathbf{0}.$ (81)
It now follows from equations (80) and (81) that condition (78) is satisfied.
Let us now focus on the condition (79). We start by calculating the projection of $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ onto the finite-dimensional tangent space $\mathcal{T}_{s_{0}}$. From (74), (71), (70) and (81), we get:
$\begin{split}\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})&=\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)=E\{\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})|\mathcal{Q}\}.\end{split}$ (82)

### v3

@@ III-A Essentials on RES distributions
$s_{s_{0}}=_{d}(2s_{0})^{-1}\left(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-m\right).$ (62)
We note in passing that the expression of $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ in (61) was already derived in [11, eq. (5)] whereas the expression for $s_{s_{0}}$ is new. Now that we have the explicit expression for the score vector $\mathbf{s}_{\bm{\eta}_{0}}$ in the model $\mathcal{P}_{\bm{\eta},\bar{g}}$ we can easily evaluate the related FIM $\mathbf{I}_{\bm{\eta}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\eta}_{0}}\mathbf{s}_{\bm{\eta}_{0}}^{T}\}$. Before computing FIM, we define two functionals of the true density generator $\bar{g}_{0}$:
$\alpha(\bar{g}_{0})\triangleq\frac{E\{\mathcal{Q}^{2}\bar{\varphi}^{2}_{0}(\mathcal{Q})\}}{m(m+2)},$ (63)
$\beta(\bar{g}_{0})\triangleq\frac{E\{\mathcal{Q}\bar{\varphi}^{2}_{0}(\mathcal{Q})\}}{m}.$ (64)
In the rest of this work, we will always assume that the density generator $\bar{g}$ is a $C^{1}$-function satisfying the regularity condition $\int_{0}^{\infty}\bar{g}(t)^{-1}(\bar{g}(t)/dt)^{2}t^{m/2+1}dt<\infty$. This guarantees that $\alpha(\bar{g}_{0})$ and $\beta(\bar{g}_{0})$ are both finite. It is important to note that this condition corresponds to the Assumption [10, Assum. 2.2] which gives sufficient conditions to assure that the elliptical distributions are a locally asymptotically normal (LAN) family.
@@
where  [^3] Note that, in the calculations, we used the equality $E\{\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\}=m$. In fact, from (58), (33) and (30), we have $E\{\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\}=-2\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2}d\bar{g}=-2\delta_{m}^{-1}[q^{m/2}\bar{g}(q)]_{0}^{\infty}+m\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2-1}\bar{g}(q)dq=m$. Similar calculation gives $E\{\mathcal{Q}^{2}\bar{\varphi}_{0}(\mathcal{Q})\}=m(m+2)$ under the additional condition $\lim_{q\rightarrow\infty}q^{m/2+1}\bar{g}(q)=0$.
$\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\}=\beta(\bar{g}_{0})[s_{0}\mathbf{V}_{S,0}]^{-1},$ (66)
@@
$\begin{split}\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}&\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}\}=\\ &={\color[rgb]{0,0,0}\frac{1}{4}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[2\alpha(\bar{g}_{0})(\mathbf{V}_{S,0}^{-1}\otimes\mathbf{V}_{S,0}^{-1})+(\alpha(\bar{g}_{0})-1){\rm vec}({\bf V}_{S,0}^{-1}){\rm vec}^{T}({\bf V}_{S,0}^{-1})][\mathbf{M}_{S}^{\mathbf{V}_{S,0}}]^{T}},\end{split}$ (69)
$\begin{split}I_{s_{0}}\triangleq E_{0}\{(s_{s_{0}})^{2}\}=\frac{m(m+2)\alpha(\bar{g}_{0})-m^{2}}{4s_{0}^{2}},\end{split}$ (70)
$\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}s_{s_{0}}\}=\frac{(m+2)\alpha(\bar{g}_{0})-m}{4s_{0}}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1}).$ (71)
We note that these expressions coincide with those already given in [11, eqs. (6) and (7)].
III-D The score vectors $\mathbf{s}_{\bm{\mu}_{0}}$, $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ in $\mathcal{P}^{u}_{\bm{\gamma}}$
Since this sub-model involves an unconstrained $g_{0}\in\mathcal{G}$, we need to define an “unconstrained” version of the function $\bar{\varphi}_{0}(t)$ in (58) as:
$\varphi_{0}(t)=s_{0}^{-1}\bar{\varphi}_{0}(s_{0}^{-1}t).$ (72)
Consequently, from (32) and from the definition of the shape matrix in (37), it is immediate to verify that the score vectors $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ in $\mathcal{P}^{u}_{\bm{\gamma}}$ are equal in distribution to the ones already derived in $\mathcal{P}^{c}_{\bm{\gamma},s}$ whose expressions are given in (56) and (61). Moreover, we have that:
$\alpha(g_{0})=\alpha(\bar{g}_{0}),\quad\beta(g_{0})=s_{0}^{-1}\beta(\bar{g}_{0}).$ (73)
IV The geometry of $\mathcal{P}_{\bm{\gamma},g}$ and $\mathcal{P}_{\bm{\eta},\bar{g}}$ and some results on parametric and semiparametric information matrices
We now state the key result that allows us to study the impact of finite and infinite dimensional nuisance parameters on the semiparametric efficiency in $\mathcal{P}_{\bm{\gamma},g}$ and $\mathcal{P}_{\bm{\eta},\bar{g}}$ for the parameter vector of interest $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$. To this end, we have the following tools at our disposal:
  1. 1.
    The score vector of $\bm{\gamma}_{0}$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ and $\mathcal{P}^{u}_{\bm{\gamma}}$
@@
    where $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ are given in (56) and (61), respectively. Moreover, we recall here that $\mathbf{s}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{H}$.
  2. 2.
    The (one-dimensional) nuisance tangent space $\mathcal{T}_{s_{0}}$ of the parametric sub-model $\mathcal{P}^{c}_{s}$ in (44) at $s_{0}$
    $\mathcal{T}_{s_{0}}=\mathrm{Span}\{s_{s_{0}}\}\subset\mathcal{H}$ (75)
    and the relevant orthogonal projection derived form (14) as:
    $\Pi(h|\mathcal{T}_{s_{0}})=E\{hs_{s_{0}}\}I_{s_{0}}^{-1}s_{s_{0}},\;h\in\mathcal{H},$ (76)
@@
  3. 3.
    The (infinite-dimensional) nuisance tangent space $\mathcal{T}^{u}_{g_{0}}$ of the unconstrained non-parametric model $\mathcal{P}^{u}_{g}$ in (49) at $g_{0}\in\mathcal{G}$:
    $\begin{split}\mathcal{T}^{u}_{g_{0}}=\left\{h\in\mathcal{H}|h\;\text{is $\sigma(\mathcal{Q})$-measurable}\right\},\end{split}$ (77)
    where $\sigma(\mathcal{Q})\subset\mathfrak{B}(\mathcal{X})$ is the sub-$\sigma$-algebra generated by the random variable $\mathcal{Q}$ in (32). A detailed derivation of this result is given in Appendix .8. From e.g. [27, Ch. 23, Def. 4], we have that the orthogonal projection of a generic element $h\in\mathcal{H}$ onto $\mathcal{T}_{g_{0}}^{u}$ can be obtained as:  [^4] It is important to note that for a generic $f\in L_{2}(P_{0})$, this projection is $\Pi(f|\mathcal{T}^{u}_{g_{0}})=E\{f|\mathcal{Q}\}-E\{f\}$.
    $\Pi(h|\mathcal{T}^{u}_{g_{0}})=E\{h|\mathcal{Q}\},\;\forall h\in\mathcal{H}.$ (78)
  4. 4.
    The (infinite-dimensional) nuisance tangent space $\mathcal{T}^{c}_{\bar{g}_{0}}$ of the constrained non-parametric model $\mathcal{P}^{c}_{\bar{g}}$ in (46) at $\bar{g}_{0}\in\overline{\mathcal{G}}$:
    $\begin{split}\mathcal{T}^{c}_{\bar{g}_{0}}=\left\{l\in\mathcal{T}^{u}_{g_{0}}|E\{(\mathcal{Q}-m)l(\mathcal{Q})\}=0\right\}=\left\{l\in\mathcal{T}^{u}_{g_{0}}|E\{\mathcal{Q}l(\mathcal{Q})\}=0\right\},\end{split}$ (79)
    and then, as a direct consequence, we have that:
    $\mathcal{H}\supseteq\mathcal{T}^{u}_{g_{0}}=\mathcal{T}^{c}_{\bar{g}_{0}}\oplus\mathrm{Span}\{\mathcal{Q}-m\}.$ (80)
    A detailed proof of this result is given in Appendix .8 along with the derivation of the relevant orthogonal projection operator that can be expressed as:
    $\Pi(h|\mathcal{T}^{c}_{\bar{g}_{0}})=E\{h|\mathcal{Q}\}-E\left\{\mathcal{Q}h\right\}\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m),\quad\forall h\in\mathcal{H}.$ (81)
    where $\sigma_{\mathcal{Q}}\triangleq\sqrt{E\left\{(\mathcal{Q}-m)^{2}\right\}}$. It is important to note here that to ensure that the decomposition in (80) and the projection operator in (81) are well-defined, we need to impose that $E\{Q^{2}\}<\infty$ (along with, of course $E\{Q\}=m$).
  5. 5.
    From its expression given in (62) and the footnote 3, we verify that, when the score $s_{s_{0}}$ has zero-mean and is $\sigma(\mathcal{Q})$-measurable, then $\mathcal{T}_{s_{0}}\subset\mathcal{T}^{u}_{g_{0}}$. However, through some simple calculation, we can verify that $E\{\mathcal{Q}s_{s_{0}}(\mathcal{Q})\}\neq 0$, then $\mathcal{T}_{s_{0}}\nsubseteq\mathcal{T}^{c}_{g_{0}}$. Finally, from [1, Theo. 5, A.4] (see also (162) and the detailed explanation in Appendix .4), we have that:
    $\Pi(h|\mathcal{T}_{s_{0}}+\mathcal{T}^{c}_{\bar{g}_{0}})=\Pi(h|\mathcal{T}^{c}_{\bar{g}_{0}})+E\{hr_{0}\}(E\{r_{0}^{2}\})^{-1}r_{0}=\Pi(h|\mathcal{T}^{u}_{g_{0}}),\;\forall h\in\mathcal{H},$ (82)
    where $r_{0}\triangleq s_{s_{0}}-\Pi(s_{s_{0}}|\mathcal{T}^{c}_{\bar{g}_{0}})$ and where the second equality holds because $r_{0}\in\mathrm{Span}\{\mathcal{Q}-m\}$.  [^5] Note that, from (62), (81) and footnote 3, we have $\Pi(s_{s_{0}}|\mathcal{T}^{c}_{\bar{g}_{0}})=E\{s_{s_{0}}|\mathcal{Q}\}-E\left\{\mathcal{Q}s_{s_{0}}\right\}\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m)=s_{s_{0}}+ms_{0}^{-1}\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m)$. Consequently, $r_{0}=ms_{0}^{-1}\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m)\in\mathrm{Span}\{\mathcal{Q}-m\}$.
We are now ready to establish an interesting result on the adaptivity property of $\bm{\gamma}_{0}$ with respect to the infinite-dimensional nuisance $g_{0}$ (or $\bar{g}_{0}$) in the presence of the finite-dimensional nuisance $s_{0}$. This surprising adaptivity result was already discovered in [11, 10] as an application to the elliptical model of the general result in [2, Prop. 2.6]. Here we will provide a more direct proof, based on the explicit form (as opposed to the asymptotic approximation used in [11, 10]) of the efficient semiparametric score vector $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$.
@@
The two efficient semiparametric FIMs $\bar{\mathbf{I}}(\bm{\gamma}_{0}|g_{0})$ for the “unconstrained” model $\mathcal{P}_{\bm{\gamma},g}$ in (39) and $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$ for the “constrained” model $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40) are both equal to the efficient parametric FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ in (45) that is a block diagonal matrix of the form
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\mathrm{blkdiag}(\mathbf{I}_{\bm{\mu}_{0}},\bar{\mathbf{I}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}),$ (83)
where $\mathbf{I}_{\bm{\mu}_{0}}$ is given in (66) and
$\bar{\mathbf{I}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=2^{-1}\alpha(\bar{g}_{0})\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[({\bf V}_{S,0}^{-1}\otimes{\bf V}_{S,0}^{-1})-m^{-1}{\rm vec}({\bf V}_{S,0}^{-1}){\rm vec}^{T}({\bf V}_{S,0}^{-1})][\mathbf{M}_{S}^{\mathbf{V}_{S,0}}]^{T}.$ (84)
@@
Let us start by proving that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$. To this end, we just need to show that $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}^{u}_{g_{0}})=\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})$. For the location parameter $\bm{\mu}_{0}$, from the expression of $\Pi(\cdot|\mathcal{T}_{s_{0}})$ in (76) and from the fact that $\mathbf{I}_{\bm{\mu}_{0},s_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}s_{s_{0}}\}=\mathbf{0}$ as indicated in (68), we immediately have that:
$\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{s_{0}})=\mathbf{I}_{\bm{\mu}_{0},s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=\mathbf{0},$ (85)
then $\mathbf{s}_{\bm{\mu}_{0}}\perp\mathcal{T}_{s_{0}}$. Moreover, from the expression of the score $\mathbf{s}_{\bm{\mu}_{0}}$ given by (56), the independence between $\mathcal{Q}$ and $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the property $E\{\mathbf{u}\}=\mathbf{0}$, we have from (78):
$\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}^{u}_{g_{0}})=E\{\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{Q}\}=\mathbf{0},$ (86)
then $\mathbf{s}_{\bm{\mu}_{0}}\perp\mathcal{T}^{u}_{g_{0}}$.
Let us now proceed to the calculation of the projection of $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ onto the finite-dimensional tangent space $\mathcal{T}_{s_{0}}$. From (76), (71), (70) and (86), we get:
$\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-m\right).$ (87)